OpenStudy (cwrw238):

I came across this problem when reading a math book. How many zeroes are at the end of 100!. One way I guess is to take out factors of 10 and see how factors of 10 arise. Any thoughts on a this?

OpenStudy (cwrw238):

100! is boviously = y * 10^x where y does not end in a zero.

OpenStudy (cwrw238):

* obviously

OpenStudy (anonymous):

Well we know 5x2 = 10, since we have more twos than fives we can just pay attention to the fives. 100/5 = 20, then 100, 75, 50, 25, all have two fives, which we already have included. = 24

OpenStudy (rational):

factors of 10 is one way

OpenStudy (rational):

the only way i think*

OpenStudy (cwrw238):

I couldn't get my head around iambatman's answer at first but I see it now. 24 is correct.

OpenStudy (anonymous):

Maybe I should've clarified by saying each set of 5 and 2 give an output of 0.

OpenStudy (anonymous):

There are 24 0's in the end of 100

OpenStudy (anonymous):

the highest power of 5 which divides 100! is given by [100/5]+[100/25] +[100/125] + [100/625]+... =20+4+0+0... =24 thus the highest power of 5 which divides 100! is 5^24.. now since there are more 2s in 100! than 5s in so 100! conatins 24 0s at the end..